A train slows from 108km/h with a uniform retardation of 5m/s2.
1.How long will it take to reach 18km/h?
2.what is the distance covered?
1
Expert's answer
2021-07-08T13:28:16-0400

"108\ km/h=30\ m/s"


"18\ km/h=5\ m/s"


(1) "v=v_0-at\to t=\frac{v_0-v}{a}=\frac{30-5}{5}=5\ (s)"


(2) "s=\frac{v^2-v_0^2}{2a}=\frac{5^2-30^2}{2\cdot(-5)}=87.5\ (m)"